find the component of a=3i+4j along the direction of i+j and i-j.here what is actually asked to  be found.which formula do we use?
 

Grade:11

1 Answers

Arun
5 years ago
Let A= 3i+4j and a(cap) be the unit vector along i+j .First find the unit vector of I+j so it will be I+j/|i+j| = 1/√2 (i+j) .Now find the Magnitude of component of 3i+4j by dot product of 3i+4j with I+j/√2 so it will be 7/√2So We got the Magnitude of Component of the vectorNow Component of 3i+2j along I+j= Magnitude of 3i+4j × unit vector of I+j=7/√2(i+j)